Find the area of the surface of circular pond that has diameter of 12m the pond is surrounded by path of width 2m if it costs $55 perm^2 to pave the path with tiles find the cost incurred

Answer by Cromlix(4381)

Hi There,
Area of circular pond = pi x radius^2 = pi x 6^2 = 113.1 m^2
Area of pond + path = pi x 8^2 = 201.1 m^2
Area of pond + path - Area of pond = 201.1 - 113.1 = 88 m^2
88 m^2 x $55 per m^2 = $4,840.
Hope this helps :-)